If Fran took 50 ml of a 25% solution how much of a 68% solution would she have to add to obtain to get a mixture of 43%
Answer by richwmiller(17219)

We have 50 mL of 25% solution
We want to add some 68% solution to make 43% solution
0.68*x+0.25*50=0.43(50+x)
0.68*x+12.5=21.5+0.43x
0.68x-0.43x=21.5-12.5
0.25x=9
x=36 mL at 68% (24.48) is added to 50 mL at 25%(12.5) to make 43% of 86 mL(36.98)
check
0.68*36+0.25*50=0.43(50+36)
24.48+12.5=0.43(86)
36.98=36.98
ok
codemixadd